# Empirical evaluation of normalizing flows in Markov Chain Monte Carlo

### v1

@@ Abstract
Keywords: normalizing flow, Markov Chain Monte Carlo, comparison, sampling, simulation study
@@ 1 Introduction
In recent years, many works have used normalizing flows (NF) within Markov Chain Monte Carlo (MCMC) to accelerate distribution sampling in lattice field theory (Del Debbio et al., 2021; Matthews et al., 2022; Abbott et al., 2023), molecular dynamics (Wu et al., 2020; Arbel et al., 2021; Matthews et al., 2022), gravitational wave analyses (Karamanis et al., 2022; Williams et al., 2021), general Bayesian model posteriors (Hoffman et al., 2019; Grumitt et al., 2022, 2024), and other fields. NF-based MCMC (NFMC) approaches typically either transform a geometrically complex target distribution into a simple one that is more amenable to MCMC sampling (Hoffman et al., 2019) or replace some MCMC steps with independent sampling from an NF, which facilitates transitions to distant parts of the sampling space (Samsonov et al., 2024). These enhancements allowed researchers to obtain satisfactory analysis results compared to classic MCMC. Still, many papers use specific combinations of MCMC samplers and NF architectures for their target distribution. A recent paper discusses some practical differences between NFMC samplers (Grenioux et al., 2023). In Section 2, we list several NFMC works and highlight key observations regarding NF properties within NFMC. Many of these propose NFMC and NF combinations for their specific problem but do not thoroughly evaluate existing methods or compare them to other NFMC and MCMC methods. The lack of systematic evaluation and comparison hinders the adoption of NFMC for sampling and leaves some key questions unanswered:
@@
The first question is crucial to understanding when to invest time in configuring the more complex NFMC samplers. The second question deals with choosing a suitable initial sampler-NF combination and saving time otherwise spent trying different, possibly unsuitable methods. Answering these questions would substantially speed up research and analysis workflows that rely on NFMC while introducing the field to a broader audience that does not necessarily have the expertise required to select appropriate sampler-NF combinations. Since NFs are largely black box models based on deep neural networks, a suitable way to answer these questions is empirically.
@@
We answer the first question by evaluating different NF architectures and NFMC samplers on various target distributions and then comparing results to MCMC. We answer the second question by narrowing our analysis to contexts where NFMC is superior to MCMC and then performing a detailed comparison of NFs within NFMC. We state which NFMC-NF combination is the best choice in general and for families of similar distributions, as well as what NF properties are correlated with good sampling performance.
@@
We investigate architectures in the family of autoregressive, residual, and continuous NFs, including popular models like Real NVP (Dinh et al., 2017), MAF (Papamakarios et al., 2017) and IAF (Kingma et al., 2016), neural spline flows (Durkan et al., 2019, NSF), invertible ResNets (Behrmann et al., 2019, i-ResNet), continuous NFs (Grathwohl et al., 2018; Salman et al., 2018), and others. We first consider Metropolis-Hastings (MH) and Hamiltonian Monte Carlo (HMC) as classic MCMC baselines when the target log density gradients are unavailable and available, respectively. We then augment each in two separate ways: by preconditioning the target according to NeuTra MCMC (Hoffman et al., 2019) and by replacing some MCMC steps with independent NF jumps according to Local-Global MCMC (Samsonov et al., 2024), which we term Jump MCMC, abbreviating the MH version as Jump MH and HMC version as Jump HMC. As a special case of Jump MH, we consider the independent Metropolis-Hastings sampler (Samsonov et al., 2024; Brofos et al., 2022, IMH), which proposes the new state at each step as a sample from an NF. We compare NFMC and NF methods on a custom benchmark consisting of four target distribution families: synthetic Gaussians, synthetic unimodal non-Gaussians, synthetic multimodal targets, and real-world distributions consisting of commonly analyzed Bayesian model posteriors. In Section 3, we describe in detail all considered NF architectures, samplers, target distributions, and our comparison methodology.
@@ 2 Related work
The most relevant work is by Grenioux et al. (2023), who analyze NeuTra MCMC, Jump MCMC, and NF-based importance sampling. They conclude that NeuTra MCMC is generally superior to Jump MCMC but does not mix between target modes. They also state that NFMC performance drops with increasing target dimensionality, with Jump MCMC being the most affected by this. However, they draw their conclusions based on only one or two synthetic targets per experiment and only four autoregressive NF architectures, which limits generalization. We build on their work by increasing the number of NF architectures with representatives from different NF families, as well as the suite of both synthetic and real-world target distributions.
@@
A recent analysis finds that IMH outperforms the Metropolis-adjusted Langevin algorithm (MALA) on a 50D Gaussian process, a 2D multimodal target, and a multimodal lattice field theory target (Brofos et al., 2022). Other works also find that IMH improves upon MCMC in field theory analyses (Albergo et al., 2019; Del Debbio et al., 2021), with Abbott et al. (2023) noting that scaling and scalability are affected by the choice of sampler hyperparameters, NF architecture, and NF hyperparameters. However, numerical experiments in these works mainly consider field theory examples and use at most a few NF architectures. Furthermore, they focus on multimodal distributions, whereas IMH could also be applied to other targets. Hoffman et al. (2019) state that sampling quality in NeuTra MCMC depends on the NF fit and that poor fits could lead to slow mixing in the tails, but only consider the IAF architecture without investigating sensitivity to hyperparameters or other architectures. Samsonov et al. (2024) show that interleaving MALA with Real NVP jumps explores a 50D multimodal target and a 128D latent space of a generative adversarial network target better than either sampler alone. Their approach outperforms the no-U-turn sampler (Hoffman and Gelman, 2011, NUTS) on the funnel and Rosenbrock distributions in low-to-moderate dimensions but performs worse in high dimensions.
@@
There are a number of findings for NF-based particle transport methods. Karamanis et al. (2022) state that their method is useful for computationally expensive targets, targets with highly correlated dimensions, and multimodal targets. However, they only consider the MAF architecture in their experiments. Grumitt et al. (2022) also apply their method to computationally expensive targets and find that small learning rates increase NF robustness for targets with complicated geometries. Grumitt et al. (2024) suggest using NF architectures with good inductive biases for common target geometries, e.g., in hierarchical Bayesian models, but do not state specific architectures or empirically explore this idea. Wu et al. (2020) find NSF to be comparable or more expressive than the commonly used Real NVP for particle transport, which suggests that NFMC methods would benefit from more careful architecture choices. They also find that their NFMC method improves upon MCMC for the double well and alanine dipeptide multimodal test cases. Arbel et al. (2021) and Matthews et al. (2022) build on this work, but both only consider the Real NVP architecture.
@@ 3 Methods
In this section, we list the analyzed MCMC methods and NF architectures. We also describe the target distributions in the benchmark and our comparison methodology.
@@ 3.1 Samplers
We first aim to show that adding NFs to MCMC can improve its performance through independent jumps or preconditioning. To eliminate some sources of variance within our experiments, we first limit ourselves to MCMC samplers with established and stable kernel tuning procedures. We focus on NeuTra MCMC and Jump MCMC as they are direct extensions of MCMC without additional sampler components. They let us assess the preconditioning and jump performance of NFs in a controlled manner. Methods like NUTS are very successful in practice. However, NUTS presently has no NeuTra MCMC or Jump MCMC extensions and theoretically developing these is beyond the scope of our paper. Furthermore, it has already been compared to NFMC in previous works (Samsonov et al., 2024; Grumitt et al., 2022). Due to the large number of both NF architectures and target distributions in the benchmark, evaluating many MCMC samplers would also lead to an even greater combinatorial explosion in the number of experiments.
We thus limit ourselves to MH as the gradient-free MCMC representative and HMC as the gradient-based one. Besides being used in various practical analyses, their NeuTra MCMC and Jump MCMC extensions have also been theoretically analyzed (Brofos et al., 2022; Samsonov et al., 2024; Hoffman et al., 2019), which facilitates the understanding and discussion of experiments in our work. Our results thus enrich the past assessments of these methods with a broader range of NF architectures. We also provide many new empirical results for Jump MH and NeuTra MH, which are not commonly used but fit exactly into the frameworks defined by Samsonov et al. (2024) and Hoffman et al. (2019). Our results may encourage their use in practical gradient-free analyses.
@@ 3.1.1 NeuTra preconditioning
For preconditioning, we consider the NeuTra method (Hoffman et al., 2019). Instead of sampling from the target density $p$, NeuTra MCMC adjusts $p$ with a bijection $f$ and samples from the adjusted density. After MCMC, we transform the sampled points with $f^{-1}$, yielding the samples from the target density. The adjusted log density is defined as:
$\log\widetilde{p}(x)=\log p(f(x))+\log\left|\det\left(\frac{\partial f(x)}{\partial x}\right)\right|.$ (1)
The bijection $f$ is associated with an NF, which is fit to $\log p$ with stochastic variational inference (SVI; see Rezende and Mohamed, 2015, for application with NFs). After SVI, the NF remains fixed and we proceed with sampling from the adjusted density. Given a fixed computational budget, the challenge is training NFs well enough to outperform MCMC despite being unable to adjust $f$ during sampling.
@@ 3.1.2 Local MCMC and global NF proposals
Another approach to using NFs within MCMC is replacing some MCMC proposals with independent sampling from an NF (Samsonov et al., 2024). We term this Jump MCMC as independent sampling with a global NF proposal amounts to jumping to different parts of space to continue MCMC exploration. The most straightforward approach is to replace the MCMC proposal with an independent sample from an NF $q$ every $K$-th iteration. Assume the chain is at iteration $t$, divisible by $K$. The proposal and log acceptance probability are:
$x^{\prime}_{t+1}$ $\sim q(x_{t}),$ (2)
$\log\alpha_{t}$ $=\log p(x^{\prime}_{t+1})-\log p(x_{t})+\log q(x_{t})-\log q(x^{\prime}_{t+1}).$ (3)
@@
When investigating NFs in these aspects, we consider two scenarios that govern which underlying MCMC samplers are sensible choices. First, we consider the case where the target log density gradient is unavailable. Such cases are found in, e.g., cosmology (Karamanis et al., 2022) or simulations of physical systems (Grumitt et al., 2024), as many existing codes for dynamical system simulations are not differentiable despite recent efforts to change this (Schoenholz and Cubuk, 2021). Second, we consider cases where the target log density has an available gradient. This covers many different Bayesian model posteriors (Agrawal and Domke, 2024; Hoffman et al., 2019) and recent work in lattice field theory (Albergo et al., 2019).
3.1.3 Other NFMC samplers
We acknowledge NFMC samplers such as Deterministic Langevin Monte Carlo (Grumitt et al., 2022, DLMC), Transport elliptical slice sampling (Cabezas and Nemeth, 2023, TESS), and Nested sampling (Williams et al., 2021, NS), however we choose to exclude them from our experiments. We justify this as follows:
@@ 3.1.3 Other NFMC samplers
  • We omit NS as it relies on an extensive sampling pipeline that requires additional hyperparameter considerations.
  • We omit TESS as a gradient-free representative, as the underlying dynamics of elliptical slice sampling are more complex and more challenging to analyze than simple MH dynamics.
We also acknowledge NFMC methods with roots in sequential Monte Carlo (Karamanis et al., 2022; Wu et al., 2020; Arbel et al., 2021; Matthews et al., 2022) and importance sampling (Midgley et al., 2023). However, the former methods rely on a more extensive set of hyperparameters (including careful target-dependent temperature scheduling), while the latter is not technically an MCMC method.
@@ 3.2 Normalizing flow architectures
An NF is a distribution defined as a transformation of a simple distribution $p_{Z}$ into a complex distribution $p_{X}$ with a bijection $f$. Sampling from an NF is equivalent to sampling $z\sim p_{Z}$ and transforming the sample with $x=f^{-1}(z)$, while $\log p_{X}$ is computed as:
$\log p_{X}(x)=\log p_{Z}(f(x))+\log\left|\det\left(\frac{\partial f(x)}{\partial x}\right)\right|.$ (4)
This expression is similar to Equation 1, which maps the target distribution to a transformed target via a bijection $f$. The bijection is most commonly parameterized by deep neural networks. The difference is in our base distribution and intended use: Equation 1 transforms the target log density into a density that is easier to sample. Equation 4 transforms a simple distribution into a complex one that acts as a global proposal distribution or whose bijection $f$ preconditions a target density. Papamakarios et al. (2021) reviewed a large number of NF architectures, all defined using Equation 4, and identified three main NF families with different approaches to constructing $f$. We describe these families in the following subsections and list the architectures we investigate.

### v2

@@ Abstract
keywords:
normalizing flow, Markov Chain Monte Carlo, comparison, sampling, simulation study
pacs:
MSC Classification]62-08
© The Author(s) 2025. This is the author’s accepted manuscript of an article accepted for publication in Machine Learning (Springer). The final authenticated version will be available at https://link.springer.com/ once published.
@@ 1 Introduction
In recent years, many works have used normalizing flows (NF) within Markov Chain Monte Carlo (MCMC) and Bayesian inference to accelerate distribution sampling in lattice field theory [16, 36, 1], molecular dynamics [51], gravitational wave analyses [32, 52], general Bayesian model posteriors [30, 24, 26], and other fields. NF-based MCMC (NFMC) approaches typically either transform a geometrically complex target distribution into a simple one that is more amenable to MCMC sampling [30] or replace some MCMC steps with independent sampling from an NF, which facilitates transitions to distant parts of the sampling space [46]. These enhancements allowed researchers to obtain satisfactory analysis results compared to classic MCMC. Still, many papers use specific combinations of MCMC samplers and NF architectures for their target distribution. A recent paper discusses some practical differences between NFMC samplers [23]. In Section 2, we list several NFMC works and highlight key observations regarding NF properties within NFMC. Many propose NFMC and NF combinations for their specific problem, but do not thoroughly evaluate existing methods or compare them to other NFMC and MCMC methods. The lack of systematic evaluation and comparison hinders the adoption of NFMC for sampling and leaves some key questions unanswered:
@@
The first question is crucial to understanding when to invest time in configuring the more complex NFMC samplers. The second question deals with choosing a suitable initial sampler-NF combination and saving time otherwise spent trying different, possibly unsuitable methods. Answering these questions would substantially speed up research and analysis workflows that rely on NFMC while introducing the field to a broader audience that does not necessarily have the expertise required to select appropriate sampler-NF combinations.
We choose to answer these questions empirically for several reasons. First, it is time-consuming to develop a reasonable theory for an NF model, a family of target distributions, or an NFMC sampler. Actionable findings can be obtained empirically and may also motivate different approaches to theoretical research. An empirical evaluation still compares the methods with useful metrics, while the comparison framework can also be used for future NF methods to quickly assess their practicality in MCMC. Lastly, we acknowledge that theoretical insights exist for training NFs in multimodal scenarios [11] and that they have been connected to NFMC [23]. Specifically, the difficulty of fitting an NF with a unimodal latent distribution increases with increasing distance between target modes. This may result in an ill-posed NF fit, where a pathological strand connects two modes in the NF distribution due to a topological mismatch between the target and the latent [11]. However, the probability of sampling points on the strand may be very small compared to the modes. In applications with global NF proposals, it may only cause a small number of additional rejections while still allowing efficient jumps between modes. Related phenomena may exist and be unaddressed by existing NFMC theory, further motivating an empirical approach.
@@
We answer the first question by evaluating different NF architectures and NFMC samplers on various target distributions and comparing results to MCMC. We answer the second question by narrowing our analysis to contexts where NFMC is superior to MCMC and then performing a detailed comparison of NFs within NFMC. We state which NFMC-NF combination is the best choice in general and for families of similar distributions, as well as what NF properties are correlated with good sampling performance.
@@
We investigate architectures in the family of autoregressive, residual, and continuous NFs, including popular models like Real NVP [19], MAF [42] and IAF [33], neural spline flows [15, NSF], invertible ResNets [8, i-ResNet], continuous NFs [22, 47], and others. We first consider Metropolis-Hastings (MH) and Hamiltonian Monte Carlo (HMC) as classic MCMC baselines when the target log density gradients are unavailable and available, respectively. We then augment each in two separate ways: by preconditioning the target according to NeuTra MCMC [30] and by replacing some MCMC steps with independent NF jumps according to Local-Global MCMC [27, 46], which we term Jump MCMC, abbreviating the MH version as Jump MH and HMC version as Jump HMC. As a special case of Jump MH, we consider the independent Metropolis-Hastings sampler [46, 7, IMH], which proposes the new state at each step as a sample from an NF. We compare NFMC and NF methods on a custom benchmark consisting of four target distribution families: synthetic Gaussians, synthetic unimodal non-Gaussians, synthetic multimodal targets, and real-world distributions consisting of commonly analyzed Bayesian model posteriors. In Section 3, we describe in detail all considered NF architectures, samplers, target distributions, and our comparison methodology.
@@ 2 Related work
The most relevant work is by [23], who analyze NeuTra MCMC, Jump MCMC, and NF-based importance sampling. They find that NeuTra MCMC tends to fail for multimodal targets, but is competitive in the unimodal case, with both claims supported by empirical and theoretical evidence. They also state that NFMC performance drops with increasing target dimensionality, with Jump MCMC being the most affected. However, they draw their conclusions based on only one or two synthetic targets per experiment and only four autoregressive NF architectures, which limits generalization. We build on their work by increasing the number of NF architectures with representatives from different NF families, as well as the suite of both synthetic and real-world target distributions.
@@
A recent analysis finds that IMH outperforms the classic Metropolis-adjusted Langevin algorithm (MALA) on a 50D Gaussian process, a 2D multimodal target, and a multimodal lattice field theory target [7]. Other works also find that IMH improves upon MCMC in field theory analyses [4, 16], with [1] noting that scaling and scalability are affected by the choice of sampler hyperparameters, NF architecture, and NF hyperparameters. However, numerical experiments in these works mainly consider field theory examples and use a few NF architectures at most. Furthermore, they focus on multimodal distributions, whereas IMH could also be applied to other targets. We note that [23] derive mixing time bounds for IMH with an isotropic Gaussian proposal, which suggest worse performance in sampling log-concave (a subset of unimodal) distributions than MCMC without NFs. We complement these findings by empirically investigating IMH performance with NFs, which do not generally attain an isotropic Gaussian fit, thus also making our experiments on unimodal targets relevant. [30] state that sampling quality in NeuTra MCMC depends on the NF fit and that poor fits could lead to slow mixing in the tails, but only consider the IAF architecture without investigating sensitivity to hyperparameters or other architectures. [46] show that interleaving MALA with Real NVP jumps explores a 50D multimodal target and a 128D latent space of a generative adversarial network target better than either sampler alone. Their approach outperforms the no-U-turn sampler [28, NUTS] on the funnel and Rosenbrock distributions in low-to-moderate dimensions but performs worse in high dimensions.
@@
There are a number of findings for general NF-based sampling methods, including particle transport methods. [32] state that their method is useful for computationally expensive targets, targets with highly correlated dimensions, and multimodal targets. However, they only consider the MAF architecture in their experiments. [24] also apply their method to computationally expensive targets and find that small learning rates increase NF robustness for targets with complicated geometries. [26] suggest using NF architectures with good inductive biases for common target geometries, e.g., in hierarchical Bayesian models, but do not state specific architectures or empirically explore this idea. [51] find NSF to be comparable or more expressive than the commonly used Real NVP for particle transport, which suggests that NFMC methods would benefit from more careful architecture choices. They also find that their NFMC method improves upon MCMC for the double well and alanine dipeptide multimodal test cases. [5] and [36] build on this work, but both only consider the Real NVP architecture. Similar to Jump MCMC, [14] recently proposed an adaptive MCMC algorithm using continuous NFs as global proposals. They train continuous NFs via a flow matching objective, which minimizes the distance between a continuous NF time-dependent vector field and a time-dependent vector field that corresponds to the target distribution. While promising, the approach relies on a temperature annealing scheme that is directly tied to NF training. This prevents a direct comparison with classic MCMC, making it difficult to isolate the quality of preconditioning or jumps, and to produce an accurate architecture comparison.
@@ 3 Methods
In this section, we describe our notation, then list the analyzed MCMC methods and NF architectures. We also describe the target distributions in the benchmark and our comparison methodology.
This section provides an overview of many methods whose conventional notation sometimes overlaps. We keep the notation similar to the referenced papers when describing each method. If two methods use the same symbol to represent different objects, we redefine the symbol in each description. This helps avoid an overwhelming number of globally defined symbols. All points, distributions, and probability density functions are $D$-dimensional unless noted otherwise. All points and samples are in $\mathbb{R}^{D}$, and all probability density functions are defined on $\mathbb{R}^{D}$ We use $X$ and $Q$ to denote the target and NF distributions. Similarly, we use $p_{X}$ and $q$ to denote the target and NF densities. We denote the partial derivative of $f$ with respect to $x$ as $\partial_{x}f(x)$. If $x$ is a vector, then $\partial_{x}f(x)$ is the Jacobian matrix of $f$. Unless noted otherwise, if $f$ is a bijection mapping between a target space and a latent space, it is understood that the forward map $f$ maps a target point to a latent point, and the inverse map $f^{-1}$ maps a latent point to a target point.
@@ 3.1 Samplers
We first aim to show that adding NFs to MCMC can improve performance through independent jumps or preconditioning. To eliminate some sources of variance within our experiments, we first limit ourselves to MCMC samplers with established and stable kernel tuning procedures. We focus on NeuTra MCMC and Jump MCMC as they are direct extensions of MCMC without additional sampler components. They let us assess the preconditioning and jump performance of NFs in a controlled manner. Methods like NUTS are very successful in practice. However, NUTS presently has no NeuTra MCMC or Jump MCMC extensions, and theoretically developing these is beyond the scope of our paper. Furthermore, it has already been compared to NFMC in previous works [46, 24]. Due to the large number of NF architectures and target distributions in the benchmark, evaluating many MCMC samplers would also lead to an even greater combinatorial explosion in the number of experiments.
We thus limit ourselves to MH as the gradient-free MCMC representative and HMC as the gradient-based one. Besides being used in various practical analyses, their NeuTra MCMC and Jump MCMC extensions have also been theoretically analyzed [7, 46, 30], which facilitates the understanding and discussion of experiments in our work. Our results thus enrich the past assessments of these methods with a broader range of NF architectures. We also provide many new empirical results for Jump MH and NeuTra MH, which are not commonly used but fit exactly into the frameworks defined by [46] and [30]. Our results may encourage their use in practical gradient-free analyses.
@@ 3.1.1 NeuTra preconditioning
For preconditioning, we consider the NeuTra method [30]. Instead of sampling from the target density $p_{X}$, NeuTra MCMC adjusts $p_{X}$ with a bijection $f$ and samples from the adjusted density. The adjusted log density is defined as:
$\log\widetilde{p}(z)=\log p_{X}(f^{-1}(z))+\log\left|\det\left(\partial_{z}f^{-1}(z)\right)\right|.$ (1)
After MCMC, we transform the sampled points $z$ with $f^{-1}(z)=x$, yielding the samples from the target density $p_{X}$. The bijection $f$ is associated with an NF, which is fit to $\log p_{X}$ with stochastic variational inference [SVI; see 43, for application with NFs]. The NF remains fixed after SVI, and we sample from the adjusted density. Given a fixed computational budget, the challenge is training NFs well enough to outperform MCMC despite being unable to adjust $f$ during sampling. [45] recently explored a similar approach based on affine transformations, consisting of a shift operation and matrix multiplication using the general linear group. While not directly tied to NF preconditioning, the work provides a framework for adaptive tuning of the preconditioner with ergodicity guarantees.
@@ 3.1.2 Local MCMC and global NF proposals
Another approach to using NFs within MCMC is replacing some MCMC proposals with independent sampling from an NF [27]. We term this Jump MCMC, as independent sampling with a global NF proposal amounts to jumping to different parts of space to continue MCMC exploration. The most straightforward approach is to replace the MCMC proposal with an independent sample from an NF $Q$ every $K$-th iteration. Let $x_{t}$ denote the chain state at iteration $t$, divisible by $K$. The proposed chain state and log acceptance probability are:
$x^{\prime}_{t+1}$ $\sim Q,$ (2)
$\log\alpha_{t}$ $=\log p_{X}(x^{\prime}_{t+1})-\log p_{X}(x_{t})+\log q(x_{t})-\log q(x^{\prime}_{t+1}).$ (3)
@@
When investigating NFs in these aspects, we consider two scenarios that govern which underlying MCMC samplers are sensible choices. First, we consider the case where the target log density gradient is unavailable. Such cases are found in, e.g., cosmology [32] or simulations of physical systems [26], as many existing codes for dynamical system simulations are not differentiable despite recent efforts to change this [44]. Second, we consider cases where the target log density has an available gradient. This covers many different Bayesian model posteriors [2, 30] and recent work in lattice field theory [4].
3.1.3 Other NFMC samplers and related methods
We acknowledge NFMC samplers such as Deterministic Langevin Monte Carlo [24, DLMC] and Transport elliptical slice sampling [13, TESS]. However, we choose to exclude them from our experiments. We justify this as follows:
@@ 3.1.3 Other NFMC samplers and related methods
  • We omit TESS as a gradient-free representative, as the underlying dynamics of elliptical slice sampling are more complex and challenging to analyze than simple MH dynamics.
Several approaches utilize NFs for preconditioning and jumps, but do not explicitly form an MCMC method. We acknowledge nested sampling with NFs [52], which performs marginal likelihood estimation by combining independent NF sampling with rejection sampling. We exclude the method from our experiments, as its primary application is marginal likelihood estimation instead of target sampling. Moreover, selecting its user-defined hyperparameters requires domain expertise and additional experimentation time to avoid high variability of results, which is beyond the scope of our analyses. We also acknowledge methods with roots in sequential Monte Carlo with NFMC mutation kernels [32, 51, 5, 36] and NF-based importance sampling [38]. However, these rely on a more extensive set of hyperparameters, including careful target-dependent temperature scheduling, which would make an accurate comparison difficult.
@@ 3.2 Normalizing flow architectures
An NF is a distribution $Q$ defined as a transformation of a simple distribution $Z$ with a bijection $f$. $f$ is typically parameterized by deep neural networks. $Z$ is typically a multivariate standard normal distribution in referenced works. We also use $Z=N(0,I)$ in this paper. Sampling $x\sim Q$ is equivalent to sampling $z\sim Z$ and transforming the sample with $x=f^{-1}(z)$. The log density $\log q$ of an NF $Q$ is computed as:
$\log q(x)=\log p_{Z}(f(x))+\log\left|\det\left(\partial_{x}f(x)\right)\right|,$ (4)
where $p_{Z}$ is the density of $Z$. This expression is similar to Equation 1. The difference is in our base distribution and intended use: Equation 1 transforms the target log density into a density that is easier to sample. Equation 4 transforms a simple distribution into a complex one that acts as a global proposal distribution or whose bijection $f$ preconditions a target density. [41] reviewed a large number of NF architectures, all defined using Equation 4, and identified three main NF families with different approaches to constructing $f$. We describe these families in the following subsections and list the architectures we investigate.